OpenStudy (anonymous):

Find the slope of the line that contains the following points. R(-3, 5), S(3, -2) Round decimal answer to nearest hundredth. What DEcimal????????????????

OpenStudy (ghazi):

-2-5/3-(-3)= -7/6=-1.16666667

OpenStudy (anonymous):

rounding to the nearest hundred would be 1.2 ?????

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

im trusting you this is an electronic quiz

OpenStudy (ghazi):

trust me :)

OpenStudy (anonymous):

ok thanks

OpenStudy (ghazi):

anytime

hero (hero):

Just use y = mx + b

OpenStudy (anonymous):

Find the slope of the line that contains the following points. X(0, 6), Y(-2, 4)

OpenStudy (anonymous):

its -2/-2 but do i simplify

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

the question doesnt say to but ....

hero (hero):

For the original question you posted: Use y = mx + b: 5 = -3m + b -2 = 3m + b Add equations to get 3 = 2b Solve for b to get 3/2 = b Then substitute b back in to get m.

hero (hero):

@777chocoholic, you shouldn't post 2 questions in the same thread

OpenStudy (ghazi):

question won't ask for the simplification of slope...just write in simplest form

OpenStudy (anonymous):

@Hero your in another world

OpenStudy (anonymous):

ok

hero (hero):

am I? You're only supposed to post one question at a time. The correct thing to do if you're done with the original question you posted is to "close" this one and post a separate question. I was showing you an alternate way of doing the other question you posted.

OpenStudy (anonymous):

canta canta canta cantaleta!!!!!!!!!!11